# RoleLLM: Benchmarking, Eliciting, and Enhancing Role-Playing Abilities of Large Language Models

### v2

@@ Abstract
The advent of Large Language Models (LLMs) has paved the way for complex tasks such as role-playing, which enhances user interactions by enabling models to imitate various characters. However, the closed-source nature of state-of-the-art LLMs and their general-purpose training limit role-playing optimization. In this paper, we introduce RoleLLM, a framework to benchmark, elicit, and enhance role-playing abilities in LLMs. RoleLLM comprises four stages: (1) Role Profile Construction for 100 roles; (2) Context-Based Instruction Generation (Context-Instruct) for role-specific knowledge extraction; (3) Role Prompting using GPT (RoleGPT) for speaking style imitation; and (4) Role-Conditioned Instruction Tuning (RoCIT) for fine-tuning open-source models along with role customization. By Context-Instruct and RoleGPT, we create RoleBench, the first systematic and fine-grained character-level benchmark dataset for role-playing with 168,093 samples. Moreover, RoCIT on RoleBench yields RoleLLaMA (English) and RoleGLM (Chinese), significantly enhancing role-playing abilities and even achieving comparable results with RoleGPT (using GPT-4)111Access models, demos, and RoleBench at https://github.com/InteractiveNLP-Team/RoleLLM-public..
@@ 1 Introduction
Large Language Models (LLMs) like ChatGPT222https://chat.openai.com/, GPT-4 (OpenAI,, 2023), and PaLM (Chowdhery et al.,, 2022) are widely considered as significant milestones in the evolution of AI. The advent of LLMs has facilitated a paradigm shift in the Natural Language Processing (NLP) community, redirecting focus from traditional downstream tasks (e.g., translation (Liu et al.,, 2020), and question-answering (Brown et al.,, 2020; Raffel et al.,, 2020)) to more complex and agent-level tasks (e.g., tool-use (Schick et al.,, 2023; Qin et al., 2023a, ), and role-playing (Shanahan et al.,, 2023)). Among these applications, role-playing aims to enable or customize LLMs to simulate various characters or personas with distinct attributes and conversational styles, which provides a more nuanced interaction experience for users, and renders LLMs more familiar, companionable and immersive (Shanahan et al.,, 2023; Wang et al., 2023b, ).
@@
However, existing open-source LLMs are predominantly trained on general domains and lack specific optimization for role-playing. Besides, while state-of-the-art (SOTA) LLMs like GPT-4 (OpenAI,, 2023) exhibit advanced role-playing capabilities, their closed-source nature imposes constraints including high API costs, unavailability of fine-tuning, and limited context window size.
@@
To mitigate these issues, several methods have been previously proposed for both closed-source and open-source models (Li et al., 2023b, ; Park et al.,, 2023; Chen et al., 2023a, ; Salemi et al.,, 2023; Wei et al.,, 2023). Nevertheless, they have the following limitations: (1) limited granularity: they mainly focus on coarse-grained personality traits, professions, or personas (Li et al., 2023b, ; Park et al.,, 2023; Wei et al.,, 2023; Chen et al., 2023a, ) (e.g., programmer, writer), neglecting more complex, finer-grained role-playing at the character level (e.g., Sherlock Holmes) for nuanced interactions and enriched experiences; (2) lack of data and benchmark: there is a lack of high-quality, diverse, and extensive open-source datasets, as well as a shortage of benchmarks for evaluation; (3) API and context costs: methods relying on closed-source models such as ChatGPT and GPT-4 (OpenAI,, 2023) cannot be freely fine-tuned and hence require all supplementary information to be included in the prompt, unnecessarily occupying the context window. Besides, API costs are prohibitively high. Therefore, exploring solutions that minimize context window utilization and are based on fine-tuning open-source models is worth researching.
@@
In this paper, we introduce RoleLLM, a role-playing framework of data construction, evaluation, and solutions for both closed-source and open-source models. In Figure 2, RoleLLM includes four key stages: (1) Role Profile Construction: we construct profiles for 95 English and 5 Chinese roles at a fine-grained character level with diverse personalities, elaborately selected from 916 English and 24 Chinese publicly available scripts; (2) Context-Based Instruction Generation (Context-Instruct): We use GPT to generate high-quality QA pairs from segmented profiles to extract role-specific knowledge; (3) Role Prompting using GPT (RoleGPT): we elicit role-playing abilities in GPT via dialogue-engineering-based role prompting, utilizing system instruction and retrieval augmentation, to generate customized responses for speaking style imitation; and (4) Role-Conditioned Instruction Tuning (RoCIT): by fine-tuning open-source LLaMA (Touvron et al.,, 2023) and ChatGLM2333https://github.com/THUDM/ChatGLM2-6B (Du et al.,, 2022; Zeng et al.,, 2022) with context-efficient role conditioning on our RoleBench, a comprehensive dataset with 168,093 role-playing samples generated by Context-Instruct and RoleGPT, we obtain RoleLLaMA and RoleGLM. Note that, to the best of our knowledge, RoleBench is the first systematic instruction-tuning dataset and benchmark for fine-grained role-playing.
@@
In our experiments, we use three Rouge-L-based metrics (Lin,, 2004) to assess models on speaking style imitation, answering accuracy, and role-specific knowledge capturing. GPT-based evaluators are also utilized, aligning with AlpacaEval (Li et al., 2023c, ). Our principal findings are: (1) dialogue engineering is favored over prompt engineering by GPT evaluators for RoleGPT; (2) RoleBench markedly improves models’ role-playing abilities, even achieving competitive results with RoleGPT in some cases; (3) RoleLLaMA exhibits robust generalization in terms of speaking style imitation and accuracy to unseen roles, requiring only role descriptions and catchphrases for effective adaptation, allowing users seamless customization of new roles; (4) system-instruction-based approach surpasses retrieval augmentation in role customization effectiveness and context efficiency; (5) Context-Instruct significantly enhances models’ knowledge about their roles, outperforming retrieval-augmentation-based methods when using noisy role profiles. We refer the readers to Appendix B for demonstrations.
@@
2 Related Work
Role-Playing. Recent advances in the LLM community have showcased the potential of LLM customization and role-playing (Wei et al.,, 2023; Shanahan et al.,, 2023; Li et al., 2023a, ; Salemi et al.,, 2023; Maas et al.,, 2023; Li et al., 2023b, ; Chen et al., 2023a, ; Park et al.,, 2023). Playing specific roles enables LLMs to become more vivid (Li et al., 2023a, ), interactive (Maas et al.,, 2023; Wang et al., 2023b, ), personalized (Salemi et al.,, 2023; Halder,, 2023), and capable of addressing complex tasks (Li et al., 2023b, ; Chen et al., 2023a, ; Qian et al.,, 2023). However, open-source LLMs still lag significantly behind SOTA closed-source LLMs such as GPT in terms of role-playing capability. Moreover, to the best of our knowledge, there exists no systematic and fine-grained benchmark to assess role-playing capability. Hence, we aim to bridge this gap by fostering the enhancement of role-playing capabilities in open-source LLMs and constructing the first benchmark for role-playing.
Data Augmentation for Instruction Tuning. Instruction tuning (aka., supervised fine-tuning, SFT) aims to enhance LLMs’ ability to follow instructions (Mishra et al.,, 2022; Longpre et al.,, 2023; Ouyang et al.,, 2022; Wang et al., 2023b, ; Zhang et al., 2023b, ). One of the main challenges for SFT is to obtain high-quality and diverse instruction data (Zhou et al.,, 2023; Wang et al., 2022a, ; Zhang et al., 2023a, ). However, manually annotating high-quality instruction data is costly. Fortunately, current SOTA LLMs such as GPT-4 (OpenAI,, 2023) are increasingly considered superior to human annotators in various tasks and domains (Gilardi et al.,, 2023; Ding et al.,, 2023), in terms of both data quality and annotation cost. Thus, employing these LLMs for data augmentation is becoming an increasingly standard practice, both for general-domain instruction-following enhancement (Wang et al., 2022a, ; Honovich et al.,, 2022; Taori et al.,, 2023; Xu et al.,, 2023; Ding et al.,, 2023) and task-specific (Qin et al., 2023b, ) or domain-specific (Wang et al., 2023a, ) adaptation. For example, Self-Instruct (Wang et al., 2022a, ) proposes expanding instruction tuning data by prompting LLMs with some seed examples. ToolLLM (Qin et al., 2023b, ) employs LLMs to construct instruction data for tool learning. However, how to augment instruction tuning data for role-playing remains under-explored. Moreover, there is still a lack of effective data augmentation methods to incorporate role-specific knowledge in synthetic data. Therefore, in our work, we explore methods for both general-domain role-playing data generation and the construction of data with role-specific knowledge.
@@ 3.1 Design Principles
3.2 RoleGPT: Eliciting Role-Playing Abilities via Dialogue Engineering
Given constraints on fine-tuning, customizing GPT for role-playing typically involves prompting, such as zero-shot custom instructions555https://openai.com/blog/custom-instructions-for-chatgpt and few-shot prompt engineering (i.e., in-context learning) (Brown et al.,, 2020; Dong et al.,, 2023), following OpenAI’s Chat Markup Language (ChatML) guidelines666https://github.com/openai/openai-python/blob/main/chatml.md. However, for ChatGPT and GPT-4 which have traded their in-context learning ability for dialogue history modeling (Fu et al.,, 2022), traditional few-shot prompt engineering is insufficient to fully elicit role-playing abilities. Thus, we modify the few-shot prompting approach to dialogue engineering, as illustrated in Box 3.2. See Appendix K.1 for detailed prompts.
Specifically, we first use the audited777The audit involves asking GPT-4 basic questions about each character, and human annotators then verify to make sure that GPT-4 knows well about the character. GPT-4 to generate character descriptions and catchphrases as the core of the custom instructions (i.e., system instructions). Then, we include an overall role-playing task instruction like “Please speak like [role_name]” and retrieve top- relevant dialogue pairs in the role profile using BM25 (Robertson and Zaragoza,, 2009) as few-shot demonstrations. By doing so, RoleGPT’s responses can capture the characters’ speaking styles and include some role-specific knowledge. However, the sparsity and noise in the profiles limit the effectiveness of knowledge discovery via retrieval augmentation.
3.3 Context-Instruct: Context-based Instruction Generation
To enhance the density of role-specific knowledge within the synthetic instruction dataset, we introduce Context-Instruct for long-text knowledge extraction and instruction data generation. The role-specific instruction data generation comprises three steps: (1) segmenting role profiles; (2) generating question-confidence-answer triplet candidates; and (3) filtering and post-processing low-quality data. We will provide a brief overview as follows. Please refer to Appendix C for details.
@@ 3.3 Context-Instruct: Context-based Instruction Generation
Given the limited context size of GPT, we meticulously partition role profiles into more manageable segments. A role profile includes (a) role description and catchphrases, as well as (b) structured dialogues (c.f., Appendix I). Segment (a) is used to obtain script-agnostic instructions, and numerous segments of (b) are used to obtain script-based instructions.
@@
As illustrated in Figure 2, in the process of generating candidates for role-specific instruction data, three elements are considered: a question () related to a given segment (i.e., context), the corresponding answer (), and a confidence score with rationale (). A LLM is used to generate these triplets for each role and segment. Our preliminary trials indicated that generating QA pairs without a confidence score resulted in lower-quality questions, often appearing incompleteness for script-based instructions due to assumptions of prior knowledge, or containing hallucinations for script-agnostic instructions due to lack of context. To address this, inspired by Lin et al., (2022) and Xiong et al., (2023), the model is prompted to also generate a confidence score with rationale to evaluate the question completeness or factualness. The prompt template includes role description, catchphrases, few-shot examples and task instructions for speaking style imitation and triplet generation888The data generated by Context-Instruct exhibits a less distinct speaking style compared to that generated by RoleGPT, due to fewer demonstrations to ensure dialogic fidelity.. The generation process yields at least 400 candidates per role with multiple model runs.
@@
The filtering procedure involves confidence-score-based filtering and de-duplication to ensure data quality and diversity. Please refer to Appendix C for more details about filtering and post-processing.
3.4 RoCIT: Role-Conditioned Instruction Tuning
@@
4 RoleBench: Benchmarking Role-Playing Abilities
In this section, we introduce how to construct RoleBench and provide an analysis. RoleBench can be used to assess role-playing capabilities and to enhance such capabilities.
@@ 4.1 Data Construction
Firstly, we meticulously select 100 representative and distinctive characters with the help of GPT-4 from a diverse range of scripts, including those from NLP Movie Scripts999https://github.com/PedroUria/NLP-Movie_Scripts, SummScreen (Chen et al.,, 2022), and manually curated Chinese scripts. Secondly, role profiles are composed of GPT-4-generated role descriptions and catchphrases, verified by authors, and structured dialogues parsed from scripts (c.f., Appendix I). Thirdly, we randomly sample 1,500 English general instructions from multiple datasets, comprising Super-NaturalInstruct (Wang et al., 2022b, ), UltraChat (Ding et al.,, 2023), and Alpaca’s (Taori et al.,, 2023). We also sample COIG (Zhang et al., 2023a, ) and BELLE’s (Yunjie et al.,, 2023) to obtain 1,479 Chinese general instructions. All sampled instructions contain no more than 100 words. We de-duplicate them based on BM25 (Robertson and Zaragoza,, 2009) similarities. Fourthly, we use RoleGPT to obtain multiple responses for each general instruction (RoleBench-general) and use Context-Instruct to generate role-specific question-answer pairs (RoleBench-specific). Lastly, the obtained raw dataset undergoes a thorough cleaning to ensure response completeness, AI and role identity concealment, and non-rejection. Please see Appendix D for further details.
@@
| # of roles | 100 |
| - # of English roles | 95 |
| - # of Chinese roles | 5 |
| # of dialogue rounds | 140,726 |
@@
| - in English | 140,225 / 1,500 |
| - in Chinese | 7,384 / 1,479 |
| - of open questions (qs) | 22,479 / 223 |
| - of commonsense qs | 37,072 / 461 |
| - of knowledge-intensive qs | 88,058 / 2,295 |
@@
| - in English | 18,949 / 18,949 |
| - in Chinese | 1,535 / 1,535 |
| - of script-agnostic qs | 13,220 / 13,220 |
| - of script-based qs | 7,164 / 7,164 |
| avg. instruction length (in words) | 25.71 |
| avg. response length (in words) | 30.48 |
@@
4.2 Data Analysis
We utilize GPT-4 API to obtain RoleBench-general-en (English) and RoleBench-general-zh (Chinese)101010https://platform.openai.com/. Context-Instruct is based on GPT-3.5 API to produce RoleBench-specific-en and RoleBench-specific-zh. Role selection, as well as description and catchphrase generation, are all executed using GPT-4 API. The parameters for GPT API calling are shown in Appendix F. Below, we provide an overview of RoleBench and refer the readers to Appendix E for the list of roles.
@@ 4.2 Data Analysis
Table 1 provides basic statistics for RoleBench. Following Wang et al., 2022a , we randomly sampled 100 instances from general and role-specific subsets respectively to evaluate the quality of RoleBench, and then we asked an expert annotator to assess their qualities in three aspects. The results in Table 2 indicate that most of them are of high quality. Please see Appendix J for valid and invalid examples.
@@
We analyze the comprehensiveness and diversity of RoleBench. Figure 4 shows verb-noun structure of RoleBench-en instructions, where the top 10 verbs (inner circle) and their top 4 direct noun objects (outer circle) are depicted, accounting for 5.6% of instructions. Besides, we categorize RoleBench roles into multiple classes and build a word cloud to show their diversity in Figure 3. The length distribution of the instructions and responses is shown in Figure 5.

### v3

@@ Abstract
The advent of Large Language Models (LLMs) has paved the way for complex tasks such as role-playing, which enhances user interactions by enabling models to imitate various characters. However, the closed-source nature of state-of-the-art LLMs and their general-purpose training limit role-playing optimization. In this paper, we introduce RoleLLM, a framework to benchmark, elicit, and enhance role-playing abilities in LLMs. RoleLLM comprises four stages: (1) Role Profile Construction for 100 roles; (2) Context-Based Instruction Generation (Context-Instruct) for role-specific knowledge extraction; (3) Role Prompting using GPT (RoleGPT) for speaking style imitation; and (4) Role-Conditioned Instruction Tuning (RoCIT) for fine-tuning open-source models along with role customization. By Context-Instruct and RoleGPT, we create RoleBench, the first systematic and fine-grained character-level benchmark dataset for role-playing with 168,093 samples. Moreover, RoCIT on RoleBench yields RoleLLaMA (English) and RoleGLM (Chinese), significantly enhancing role-playing abilities and even achieving comparable results with RoleGPT (using GPT-4)111All resources are available in https://github.com/InteractiveNLP-Team/RoleLLM-public.
RoleLLM: Benchmarking, Eliciting, and Enhancing Role-Playing Abilities of Large Language Models
Zekun Moore Wang1,9∗, Zhongyuan Peng2∗, Haoran Que1∗, Jiaheng Liu1†, Wangchunshu Zhou3, Yuhan Wu8, Hongcheng Guo1, Ruitong Gan4, Zehao Ni2, Jian Yang1, Man Zhang8, Zhaoxiang Zhang5†, Wanli Ouyang6, Ke Xu1, Stephen W. Huang7, Jie Fu9, Junran Peng5 1Beihang University; 2University of the Chinese Academy of Sciences; 3ETH Zürich; 4The Hong Kong Polytechnic University; 5Institute of Automation, Chinese Academy of Sciences; 6Shanghai AI Lab; 7Harmony.AI; 8Beijing University of Posts and Telecommunications; 9The Hong Kong University of Science and Technology; zenmoore@buaa.edu.cn, liujiaheng@buaa.edu.cn
@@ 1 Introduction
Large Language Models (LLMs) like ChatGPT222https://chat.openai.com/, GPT-4 (OpenAI, 2023), and PaLM (Chowdhery et al., 2022) are widely considered as significant milestones in the evolution of AI. The advent of LLMs has facilitated a paradigm shift in the Natural Language Processing (NLP) community, redirecting focus from traditional downstream tasks (e.g., translation (Liu et al., 2020), and question-answering (Brown et al., 2020; Raffel et al., 2020)) to more complex and agent-level tasks (e.g., tool-use (Schick et al., 2023; Qin et al., 2023a), and role-playing (Shanahan et al., 2023)). Among these applications, role-playing aims to enable or customize LLMs to simulate various characters or personas with distinct attributes and conversational styles, which provides a more nuanced interaction experience for users, and renders LLMs more familiar and companionable (Wang et al., 2023b).
@@
However, existing open-source LLMs are predominantly trained on general domains and lack specific optimization for role-playing. Besides, while state-of-the-art (SOTA) LLMs like GPT-4 (OpenAI, 2023) exhibit advanced role-playing capabilities, their closed-source nature imposes constraints including high API costs, unavailability of fine-tuning, and limited context window size.
@@
To mitigate these issues, several methods have been previously proposed for both closed-source and open-source models (Li et al., 2023b; Park et al., 2023; Chen et al., 2023a; Salemi et al., 2023; Wei et al., 2023). Nevertheless, they have the following limitations: (1) limited granularity: they mainly focus on coarse-grained personality traits, professions, or personas (Li et al., 2023b; Park et al., 2023; Wei et al., 2023; Chen et al., 2023a) (e.g., programmer, writer), neglecting more complex, finer-grained role-playing at the character level (e.g., Sherlock Holmes) for nuanced interactions and enriched experiences; (2) lack of data and benchmark: there is a lack of high-quality, diverse, and extensive open-source datasets, as well as a shortage of benchmarks for evaluation; (3) API and context costs: methods relying on closed-source models such as ChatGPT and GPT-4 (OpenAI, 2023) cannot be freely fine-tuned and hence require all supplementary information to be included in the prompt, unnecessarily occupying the context window. Besides, API costs are prohibitively high. Therefore, exploring solutions that minimize context window utilization and are based on fine-tuning open-source models is worth researching.
@@
In this paper, we introduce RoleLLM, a role-playing framework of data construction, evaluation, and solutions for both closed-source and open-source models. In Figure 1, RoleLLM includes four key stages: (1) Role Profile Construction: we construct profiles for 95 English and 5 Chinese roles at a fine-grained character level with diverse personalities, selected from 916 English and 24 Chinese scripts; (2) Context-Based Instruction Generation (Context-Instruct): we use GPT to generate high-quality QA pairs from segmented profiles to extract role-specific knowledge; (3) Role Prompting using GPT (RoleGPT): we elicit role-playing abilities in GPT via dialogue-engineering-based role prompting, utilizing system instruction and retrieval augmentation, to generate responses for speaking style imitation; and (4) Role-Conditioned Instruction Tuning (RoCIT): by fine-tuning open-source LLaMA (Touvron et al., 2023) and ChatGLM2333https://github.com/THUDM/ChatGLM2-6B (Du et al., 2022; Zeng et al., 2022) with context-efficient role conditioning on RoleBench with 168,093 role-playing samples generated by Context-Instruct and RoleGPT, we obtain RoleLLaMA and RoleGLM. Note that, to the best of our knowledge, RoleBench is the first systematic instruction-tuning dataset and benchmark for fine-grained role-playing.
@@
In our experiments, we use three Rouge-L metrics (Lin, 2004), GPT, and human evaluation to assess models on speaking style imitation, answering accuracy, and role-specific knowledge capturing. Our principal findings are: (1) dialogue engineering is favored over prompt engineering by GPT evaluators for RoleGPT; (2) RoleBench markedly improves models’ role-playing abilities, achieving competitive results with RoleGPT in some cases; (3) RoleLLaMA exhibits robust generalization in terms of speaking style imitation and accuracy to unseen roles, requiring only role descriptions and catchphrases for effective adaptation, allowing users seamless customization of new roles; (4) system-instruction-based approach surpasses retrieval augmentation in role customization effectiveness and context efficiency; (5) Context-Instruct significantly enhances models’ knowledge about their roles, outperforming retrieval-augmentation-based methods when using noisy role profiles. We refer the readers to Appendix A for demonstrations and comparative case studies.
@@ 2.1 Design Principles
2.2 RoleGPT
Given constraints on fine-tuning, customizing ChatGPT for role-playing typically involves prompting, such as zero-shot custom instructions555https://openai.com/blog/custom-instructions-for-chatgpt and few-shot prompt engineering (i.e., in-context learning) (Brown et al., 2020; Dong et al., 2023). However, for ChatGPT and GPT-4 which have traded their in-context learning ability for dialogue history modeling (Fu et al., 2022), traditional few-shot prompt engineering is insufficient to fully elicit role-playing abilities. Thus, we modify the few-shot prompting approach to dialogue engineering, as illustrated in Box 2.2 (Please see Appendix I.1 for detailed prompts).
Specifically, we first use the audited666The audit involves asking GPT-4 basic questions about each character, and human annotators then verify to make sure that GPT-4 knows well about the character. GPT-4 to generate character descriptions and catchphrases as the core of the custom instructions (i.e., system instructions). Then, we include an overall role-playing task instruction like “Please speak like [role_name]” and retrieve top- relevant dialogue pairs in the role profile using BM25 (Robertson and Zaragoza, 2009) as few-shot demonstrations. By doing so, RoleGPT’s responses can capture the characters’ speaking styles and include some role-specific knowledge. However, the sparsity and noise in the profiles limit the effectiveness of knowledge discovery via retrieval augmentation.
2.3 Context-Instruct
To enhance the density of role-specific knowledge within the synthetic instruction dataset, we introduce Context-Instruct for long-text knowledge extraction and instruction generation. The role-specific instruction generation comprises three steps: (1) segmenting role profiles; (2) generating question-confidence-answer triplet candidates; and (3) filtering and post-processing low-quality data. We will provide a brief overview as follows (Please refer to Appendix B for details).
@@ 2.3 Context-Instruct
Given the limited context size of GPT, we meticulously partition role profiles into more manageable segments. A role profile includes (a) role description and catchphrases, as well as (b) structured dialogues (c.f., Appendix H). Segment (a) is used to obtain script-agnostic instructions, and numerous segments of (b) are used to obtain script-based instructions.
@@
As illustrated in Figure 1, in the process of generating candidates for role-specific instruction data, three elements are considered: a question () related to a given segment (i.e., context), the corresponding answer (), and a confidence score with rationale (). A LLM is used to generate these triplets for each role and segment. We observed that generating QA pairs without a confidence score resulted in lower-quality questions, often appearing incompleteness for script-based instructions due to assumptions of prior knowledge, or containing hallucinations for script-agnostic instructions due to lack of context. To address this, inspired by Lin et al. (2022) and Xiong et al. (2023), the model is prompted to also generate a confidence score with rationale to evaluate the question completeness or factualness. The prompt template includes role description, catchphrases, few-shot examples and task instructions for speaking style imitation and triplet generation777The data generated by Context-Instruct exhibits a less distinct speaking style compared to that generated by RoleGPT, due to fewer demonstrations to ensure dialogic fidelity.. The generation process yields at least 400 candidates per role with multiple model runs.
@@
The filtering procedure involves confidence-score-based filtering and de-duplication to ensure data quality and diversity. Please refer to Appendix B for more details about filtering and post-processing.
2.4 RoCIT
@@
3 RoleBench
In this section, we introduce the details of RoleBench, which can be used to assess and enhance role-playing capabilities.
@@ 3.1 Data Construction
Firstly, we meticulously select 100 representative and distinctive characters with the help of GPT-4 from a diverse range of scripts, including those from NLP Movie Scripts999https://github.com/PedroUria/NLP-Movie_Scripts, SummScreen (Chen et al., 2022), and manually curated Chinese scripts. Secondly, role profiles are composed of GPT-4-generated role descriptions and catchphrases, verified by authors, and structured dialogues parsed from scripts (c.f., Appendix H). Thirdly, we randomly sample 1,500 English general instructions from multiple datasets, comprising Super-NaturalInstruct (Wang et al., 2022b), UltraChat (Ding et al., 2023), and Alpaca’s (Taori et al., 2023). We also sample COIG (Zhang et al., 2023a) and BELLE’s (Yunjie et al., 2023) to obtain 1,479 Chinese general instructions. All sampled instructions contain no more than 100 words. We de-duplicate them based on BM25 (Robertson and Zaragoza, 2009) similarities. Fourthly, we use RoleGPT to obtain multiple responses for each general instruction and use Context-Instruct to generate role-specific question-answer pairs. Lastly, the obtained raw dataset undergoes a thorough cleaning to ensure response completeness, AI and role identity concealment, and non-rejection. Please see Appendix C for further details.
3.2 Data Analysis
We utilize GPT-4 API to obtain RoleBench-general-en (English) and RoleBench-general-zh (Chinese)101010https://platform.openai.com/. Context-Instruct is based on GPT-3.5 API to produce RoleBench-specific-en and RoleBench-specific-zh. Role selection, as well as description and catchphrase generation, are all executed using GPT-4 API. The parameters for GPT API calling are shown in Appendix E. Below, we provide an overview of RoleBench and refer the readers to Appendix D.3 for the list of roles.
@@ 3.2 Data Analysis
| # of English roles | 95 |
| # of Chinese roles | 5 |
@@
Table 1 provides basic statistics for RoleBench. Following Wang et al. (2022a), we randomly sampled 100 instances from general and role-specific subsets respectively to evaluate the quality of RoleBench, and then we asked expert annotators to assess their qualities in three aspects. The results in Table 2 indicate that most of them are of high quality. Please see Appendix J for valid and invalid examples.
@@
We analyze diversity of RoleBench. Figure 4 shows verb-noun structure of RoleBench-en instructions, where the top 10 verbs (inner circle) and their top 4 direct noun objects (outer circle) are depicted with 5.6% of instructions. Besides, we categorize RoleBench roles into multiple classes and build a word cloud to show their diversity in Figure 3. The length distributions of instructions and responses are shown in Figure 5.
Example.
Box 3.2 is an example in RoleBench:
Please refer to Appendix J for more examples.